In a normal distribution 41% of items are under 55 and 11% are over 66. Find the mean and variance of the distribution.


1
Expert's answer
2022-03-03T11:54:00-0500
"P(Z<\dfrac{55-\mu}{\sigma})=0.41"

"\dfrac{55-\mu}{\sigma}\approx-0.227545"



"P(Z>\dfrac{66-\mu}{\sigma})=0.11"

"\dfrac{66-\mu}{\sigma}\approx1.226528"


"\mu=66-1.226528\sigma"

"66-1.226528\sigma=55+0.227545\sigma"

"\sigma=7.565"

"\mu=56.72"